OpenStudy (anonymous):

A man starts his day with 79 bills totaling $4000 in all 100, 50, and 20 dollar bills. For lunch, he buys himself and friend lunch and tips generously. His bill comes to $2550. To pay for lunch, he uses all of his 100 dollar bills and half of his 50 dollar bills. Using this information, how many bills of each denomination did he start his day with?

OpenStudy (anonymous):

that is a hell of a lunch !

OpenStudy (anonymous):

good question too!

OpenStudy (anonymous):

Must have been quite good! so do you have any ideas on this one?

OpenStudy (anonymous):

Got it, 100x+50y+20z=4000 100x+25y+0z=2550 x+y+z=79 solved as a matrix, reduced row echelon form gives x,y,z=16,38,25 16 100s, 38 50's and 25 20's